OpenStudy (anonymous):

Find the general solution of the differential equation. Use C for the constant of integration. dy/dx= 4x + 3/x

OpenStudy (unklerhaukus):

\[\frac{\mathrm dy}{\mathrm dx}=4x+3/x\] separate the variables \[\mathrm dy=(4x+3/x)\,\mathrm dx\] now integrate both sides \[\int\mathrm dy=\int(4x+3/x)\,\mathrm dx\]

OpenStudy (anonymous):

I had 2x^2+ 3ln(x) +C and idk why its wrong

OpenStudy (unklerhaukus):

that's right \[y(x) = 2x^2+3\ln x+C\]

OpenStudy (anonymous):

but my homework says its wrong ;/ does the lnx have to be in absolute value

OpenStudy (unklerhaukus):

oh yeah, we can't take the log of a negative number, so the argument of the log should be an absolute value \[y(x) = 2x^2+3\ln|x|+C\]

OpenStudy (anonymous):

okay hang on let me see if it takes that!

OpenStudy (anonymous):

can you help me with this one? Find the general solution of the differential equation. Use C for the constant of integration. dr/dp = 10sin(p)

OpenStudy (anonymous):

the previous one was right by the way! thanks :)

OpenStudy (unklerhaukus):

Oh good. This one is the same idea, separate the variables, and integrate

OpenStudy (anonymous):

so it'd be 10xcos(p)+C..?

OpenStudy (unklerhaukus):

careful with your sign,

OpenStudy (unklerhaukus):

\[\int\sin=-\cos\]

OpenStudy (unklerhaukus):

@anarvizo